# Supplementary Material For Self-supervised Likelihood Estimation with Energy Guidance for Anomaly Segmentation in Urban Scenes

### v3

@@
Results on Lost & Found. To make more comprehensive comparison, we also evaluate our method on the old Lost & Found dataset (lostAndFound2016), where 13 challenging real-world scenes are included with 37 different anomalous instances, which is the first publicly available urban anomaly segmentation datasets. Specifically, the obstacles in this dataset vary significantly in size and material. And following the official protocol (lostAndFound2016), 1,203 images collected from 112 video stereo sequences with a resolution of 2048$\times$1024 are utilized as the test set. Similar to the setting on FS Lost & Found, the segmentation model of DeepLab series (chen2018encoder) with ResNet101 pretrained on Cityscapes is used and fixed without re-training on this dataset. Finally, the same OoD head and evaluation metrics are adopted as FS Lost & Found.
[…]
Meta-OoD (chan2021entropy) 97.95 71.23 5.95
${}^{\dagger}$ (di2021pixel) 98.38 70.43 4.89
Deep ${}^{\dagger}$ (liu2019deep) 98.67 72.73 3.81
${}^{\dagger}$ (di2021pixel) 99.76 78.29 0.81
${}^{\dagger}$ (grcic2022densehybrid) 99.37 78.70 2.10
MSP (hendrycks2019scaling) 85.49 38.20 18.56
Mahalanobis (lee2018simple) 79.53 42.56 24.51
Max Logit (hendrycks2016baseline) 94.52 65.45 15.56
Entropy (hendrycks2016baseline) 86.52 50.66 16.95
${}^{\dagger}$ (jung2021standardized) 88.05 25.89 44.48
Energy (liu2020energy) 94.45 66.37 15.69
@@
Tab. 2 shows the performance of the proposed SLEEG on the test set of Lost & Found. Notably, our method surpasses all previous approaches and achieves SOTA performance on AP. Specifically, though previous SOTA methods DenseHybrid (grcic2022densehybrid) and PEBAL (tian2021pixel) utilize auxiliary data and are re-trained with more complex WideResNet38 as backbone, our SLEEG still brings a relative improvement of 4.59% and 4.18% to them on AP respectively with ResNet101 as backbone and requires no auxiliary OoD samples or further re-training of the segmentation model. Furthermore, when compared with the SOTA method Energy (liu2020energy) that falls into the same category as SLEEG, our method surpasses it by a large gap of 16.51% on AP and meanwhile significantly reduces the false positive pixels from 15.69% to 4.6%. This results fully illustrate that SLEEG shows great robustness and effectiveness on localising various real-world unexpected objects. Moreover, SML (jung2021standardized) attempts to tackle this task by re-balance the class-wise discrepancy of inlier samples. However, since there are only two classes in Lost & Found, there exists large performance gap between the performance of SML on Fishyscapes and Lost & Found, implying that SML may not be suitable for real-world applications. By contrast, it is worth noting that our SLEEG is capable of achieving consistent performance boost on all these datasets, demonstrating its great potential in real world scenarios.
@@
DenseHybrid + OoD data  (grcic2022densehybrid) 6.1 63.8 -
@@
Table 3: Investigation when combining other state-of-the-art method DenseHybrid (grcic2022densehybrid) with our strategy of patch generation on FS LAF validation set.
Table 4: Performance evaluation on the SMIYC benchmark (chan21arxiv).
@@
Data training $\mathrm{FPR}_{95}$downarrow AP $\uparrow$
SynBoost (biase21cvpr) 61.9 56.4
JSRNet (vojir21iccv) 43.9 33.6
Void Classifier (blum21ijcv) 63.5 36.6
DenseHybrid (grcic2022densehybrid) 9.8 78.0
PEBAL (tian2021pixel) 40.8 49.1
Image Resyn. (lis19iccv) 25.9 52.3
Embed. Dens. (blum21ijcv) 70.8 37.5
ODIN (liang18iclr) 71.7 33.1
MC Dropout (kendall17nips) 69.5 28.9
Max softmax (hendrycks17iclr) 72.1 28.0
Mahalanobis (lee18nips) 87.0 20.0
@@
Table 5: Performance evaluation on StreetHazards (streethazards2019). SLEEG achieves competitive anomaly detection performance.
@@
Energy (liu20neurips) 18.2 93.0 12.9
Outlier Exposure (hendrycks19iclr) 17.7 94.0 14.6
OOD-Head (bevandic19gcpr) 56.2 88.8 19.7
OH-MSP (bevandic21arxiv) 30.9 89.7 18.8
DenseHybrid (grcic2022densehybrid) 13.0 95.6 30.2
CoroCL (liu2023residual) 8.2 97.2 31.2
SynthCP (xia20eccv) 28.4 88.5 9.3
Dropout (kendall17nips)(xia20eccv) 79.4 69.9 7.5
TRADI (franchi20eccv) 25.3 89.2 7.2
OVNNI (franchi20arxiv) 22.2 91.2 12.6
SO+H (grcic21visapp) 25.2 91.7 12.7
DML (cen21iccv) 17.3 93.7 14.7
MSP (hendrycks17iclr) 27.9 90.1 7.5
ML (hendrycks19arxiv) 22.5 92.4 11.6
ODIN (liang18iclr) 28.7 90.0 7.0
ReAct (sun21neurips) 21.2 92.3 10.9
@@
Results for OoD data replacement. Meanwhile, we also separately investigate our patch policy with other method (grcic2022densehybrid) by replacing the external OoD data used in original paper with our generated patches. The results can be illustrated in Tab. 3, where the results of Densehybrid+OoD data is borrowed from the original paper, and the other results are reimplemented by ourselves. It can be observed that although applying less data, our patch policy can achieve results close to those in the original paper, this further demonstrate that our scheme can be viewed as a potential replacement for OoD data relied by previous methods.
@@
Fig. 6 further visualizes the comparison between JEM and our SLEEG on Lost & Found test set. Notably, SLEEG can consistently perform better than JEM when detecting objects with various scales and distances. Beside JEM, Additionally, Fig. 6 and Fig. 5 describe the comparison of visualization results between the classic softmax entropy (MSP) and image re-synthesis (imgr) methods and our SLEEG. It can be observed that these methods perform much worse than our SLEEG and are unable to effectively distinguish in-distribution objects from anomalous ones. And there even exists large background regions that are assigned with higher values than anomalous objects. By contrast, our SLEEG generates anomaly maps where unexpected instances are intensively assigned with higher scores, resulting in much fewer false positive pixels as well.
@@
As shown in the Table 1 of the manuscript and Fig. 3, our SLEEG performs inferior to SOTA methods that utilize auxiliary OoD data on the FS Static dataset. This mainly accounts for that the synthetic anomalous objects in FS Static are similar to the instances in the auxiliary dataset that these methods use, i.e., COCO (lin2014microsoft) and ADE20K (zhou2019semantic). Therefore, training with such OoD samples allows models foresee the anomalous objects, resulting in their good performance. Moreover, training in this manner also results in a large gap between these methods and ours in performance on the FS Lost & Found and Lost & Found datasets since anomalous objects in the FS Lost & Found dataset are more realistic. Finally, though our SLEEG performs inferior on FS Static dataset, since the anomaly segmentation task targets at real-world automatic driving scenarios, the ability of tackling realistic unexpected samples is more essential for our SLEEG. Therefore, we focus on utilizing the inherent spatial context to improve the ability of tackling with the distribution discrepancy of normal and anomalies without requiring extra OoD data.

### v4

# Self-Supervised Likelihood Estimation with Energy Guidance for Anomaly Segmentation in Urban Scenes

@@ Acknowledgement
Results on Lost & Found. To make more comprehensive comparison, we also evaluate our method on the old Lost & Found dataset (Pinggera et al. 2016), where 13 challenging real-world scenes are included with 37 different anomalous instances, which is the first publicly available urban anomaly segmentation datasets. Specifically, the obstacles in this dataset vary significantly in size and material. And following the official protocol (Pinggera et al. 2016), 1,203 images collected from 112 video stereo sequences with a resolution of 2048$\times$1024 are utilized as the test set. Similar to the setting on FS Lost & Found, the segmentation model of DeepLab series (Chen et al. 2018) with ResNet101 pretrained on Cityscapes is used and fixed without re-training on this dataset. Finally, the same OoD head and evaluation metrics are adopted as FS Lost & Found.
Methods OoD Data Re-training AUC $\uparrow$ AP $\uparrow$ FPR95 $\downarrow$
Meta-OoD (Chan, Rottmann, and Gottschalk 2021) 97.95 71.23 5.95
SynBoost (Di Biase et al. 2021b) 98.38 70.43 4.89
Deep Gambler (Liu et al. 2019) 98.67 72.73 3.81
PEBAL (Di Biase et al. 2021b) 99.76 78.29 0.81
DenseHybrid (Grcić, Bevandić, and Šegvić 2022) 99.37 78.70 2.10
MSP (Hendrycks et al. 2019c) 85.49 38.20 18.56
Mahalanobis (Lee et al. 2018a) 79.53 42.56 24.51
Max Logit (Hendrycks and Gimpel 2016) 94.52 65.45 15.56
Entropy (Hendrycks and Gimpel 2016) 86.52 50.66 16.95
SML (Jung et al. 2021) 88.05 25.89 44.48
Energy (Liu et al. 2020a) 94.45 66.37 15.69
@@
Tab. 8 shows the performance of the proposed SLEEG on the test set of Lost & Found. Notably, our method surpasses all previous approaches and achieves SOTA performance on AP. Specifically, though previous SOTA methods DenseHybrid (Grcić, Bevandić, and Šegvić 2022) and PEBAL (Tian et al. 2021) utilize auxiliary data and are re-trained with more complex WideResNet38 as backbone, our SLEEG still brings a relative improvement of 4.59% and 4.18% to them on AP respectively with ResNet101 as backbone and requires no auxiliary OoD samples or further re-training of the segmentation model. Furthermore, when compared with the SOTA method Energy (Liu et al. 2020a) that falls into the same category as SLEEG, our method surpasses it by a large gap of 16.51% on AP and meanwhile significantly reduces the false positive pixels from 15.69% to 4.6%. This results fully illustrate that SLEEG shows great robustness and effectiveness on localising various real-world unexpected objects. Moreover, SML (Jung et al. 2021) attempts to tackle this task by re-balance the class-wise discrepancy of inlier samples. However, since there are only two classes in Lost & Found, there exists large performance gap between the performance of SML on Fishyscapes and Lost & Found, implying that SML may not be suitable for real-world applications. By contrast, it is worth noting that our SLEEG is capable of achieving consistent performance boost on all these datasets, demonstrating its great potential in real world scenarios.
@@
DenseHybrid + OoD data  (Grcić, Bevandić, and Šegvić 2022) 6.1 63.8 -
@@
Table 9: Investigation when combining other state-of-the-art method DenseHybrid (Grcić, Bevandić, and Šegvić 2022) with our strategy of patch generation on FS LAF validation set.
Table 10: Performance evaluation on the SMIYC benchmark (Chan et al. 2021b).
@@
Data training $\mathrm{FPR}_{95}$    $\downarrow$ AP $\uparrow$
SynBoost (Biase et al. 2021) 61.9 56.4
JSRNet (Vojir et al. 2021b) 43.9 33.6
Void Classifier (Blum et al. 2021b) 63.5 36.6
DenseHybrid (Grcić, Bevandić, and Šegvić 2022) 9.8 78.0
PEBAL (Tian et al. 2021) 40.8 49.1
Image Resyn. (Lis et al. 2019c) 25.9 52.3
Embed. Dens. (Blum et al. 2021b) 70.8 37.5
ODIN (Liang, Li, and Srikant ) 71.7 33.1
MC Dropout (Kendall and Gal 2017) 69.5 28.9
Max softmax (Hendrycks and Gimpel ) 72.1 28.0
Mahalanobis (Lee et al. 2018b) 87.0 20.0
@@
Table 11: Performance evaluation on StreetHazards (Hendrycks et al. 2019a). SLEEG achieves competitive anomaly detection performance.
@@
Energy (Liu et al. 2020b) 18.2 93.0 12.9
Outlier Exposure (Hendrycks, Mazeika, and Dietterich ) 17.7 94.0 14.6
OOD-Head (Bevandic et al. ) 56.2 88.8 19.7
OH-MSP (Bevandic et al. 2021) 30.9 89.7 18.8
DenseHybrid (Grcić, Bevandić, and Šegvić 2022) 13.0 95.6 30.2
CoroCL (Liu et al. 2023) 8.2 97.2 31.2
SynthCP (Xia et al. 2020b) 28.4 88.5 9.3
Dropout (Kendall and Gal 2017)(Xia et al. 2020b) 79.4 69.9 7.5
TRADI (Franchi et al. 2020b) 25.3 89.2 7.2
OVNNI (Franchi et al. 2020a) 22.2 91.2 12.6
SO+H (Grcić, Bevandić, and Šegvić ) 25.2 91.7 12.7
DML (Cen et al. 2021) 17.3 93.7 14.7
MSP (Hendrycks and Gimpel ) 27.9 90.1 7.5
ML (Hendrycks et al. 2019d) 22.5 92.4 11.6
ODIN (Liang, Li, and Srikant ) 28.7 90.0 7.0
ReAct (Sun, Guo, and Li 2021) 21.2 92.3 10.9
@@
Results for OoD data replacement. Meanwhile, we also separately investigate our patch policy with other method (Grcić, Bevandić, and Šegvić 2022) by replacing the external OoD data used in original paper with our generated patches. The results can be illustrated in Tab. 9, where the results of Densehybrid+OoD data is borrowed from the original paper, and the other results are reimplemented by ourselves. It can be observed that although applying less data, our patch policy can achieve results close to those in the original paper, this further demonstrate that our scheme can be viewed as a potential replacement for OoD data relied by previous methods.
@@
Fig. 10 further visualizes the comparison between JEM and our SLEEG on Lost & Found test set. Notably, SLEEG can consistently perform better than JEM when detecting objects with various scales and distances. Beside JEM, Additionally, Fig. 10 and Fig. 9 describe the comparison of visualization results between the classic softmax entropy (Hendrycks and Gimpel 2017) and image re-synthesis (Lis et al. 2019a) methods and our SLEEG. It can be observed that these methods perform much worse than our SLEEG and are unable to effectively distinguish in-distribution objects from anomalous ones. And there even exists large background regions that are assigned with higher values than anomalous objects. By contrast, our SLEEG generates anomaly maps where unexpected instances are intensively assigned with higher scores, resulting in much fewer false positive pixels as well.
@@
As shown in the Table 1 of the manuscript and Fig. 7, our SLEEG performs inferior to SOTA methods that utilize auxiliary OoD data on the FS Static dataset. This mainly accounts for that the synthetic anomalous objects in FS Static are similar to the instances in the auxiliary dataset that these methods use, i.e., COCO (Lin et al. 2014) and ADE20K (Zhou et al. 2019). Therefore, training with such OoD samples allows models foresee the anomalous objects, resulting in their good performance. Moreover, training in this manner also results in a large gap between these methods and ours in performance on the FS Lost & Found and Lost & Found datasets since anomalous objects in the FS Lost & Found dataset are more realistic. Finally, though our SLEEG performs inferior on FS Static dataset, since the anomaly segmentation task targets at real-world automatic driving scenarios, the ability of tackling realistic unexpected samples is more essential for our SLEEG. Therefore, we focus on utilizing the inherent spatial context to improve the ability of tackling with the distribution discrepancy of normal and anomalies without requiring extra OoD data.